# arXiv:2603.00422v1[stat.AP] 28 Feb 2026

### 28 Feb 2026 v1

@@ Abstract
Tourism demand forecasting is methodologically mature, but it typically treats accommodation supply as fixed or exogenous. In platform-mediated short-term rentals, supply is elastic, decision-driven, and co-evolves with demand through pricing, information design, and interventions. I reframe the core issue as endogenous stock-out censoring: realized booked nights satisfy $B_{k,t}\leq\min(D_{k,t},S_{k,t})$, so booking models that ignore supply learn a regime-specific ceiling and become fragile under policy changes and supply shocks. This narrated review synthesizes work from tourism forecasting, revenue management, two-sided market economics, and Bayesian time-series methods; develops a three-part coupling framework (behavioral, informational, intervention); and illustrates the identification failure with a toy simulation. I conclude with a focused research agenda for jointly forecasting supply, demand, and their compositions.
@@ 1 Introduction
The crucial shift is that “capacity” is no longer slow-moving and near-fixed (as in most hotel markets), but elastic at multiple margins: host entry and exit, listing activation and deactivation, and day-level availability decisions. Empirically, peer-to-peer platforms exhibit greater supply-side flexibility than hotels, with effective capacity that can expand or contract at the margins of host entry, listing activation, and calendar availability; the largest competitive effects concentrate in peak-demand periods when this incremental flexibility matters most (Farronato and Fradkin, 2022; Zervas et al., 2017). From a theory standpoint, two-sided market economics provides the backbone for why forecasting demand alone is incomplete in such settings: platforms internalize cross-side externalities via pricing, matching, governance, and information design (Rochet and Tirole, 2003; Weyl, 2010; Armstrong, 2006).
@@
The paper makes three contributions. First, it develops a coupling framework (behavioral, informational, and intervention coupling) that explains why independent supply and demand forecasts are fragile under policy changes, supply regime shifts, and platform intervention, even when they may perform adequately for short-horizon extrapolation in stable environments. Second, it demonstrates the resulting identification failure through a simulation showing that demand-only forecasters systematically over-predict after supply shocks while coupled forecasters do not. Third, it outlines a five-problem research agenda whose resolution would constitute the core of a coupled forecasting research program for platform markets. A detailed survey of the methodological building blocks for coupled forecasting is provided in the Supplementary Material.
@@
Let $D_{t}$ denote latent demand intent: the number of listing-nights that guests would book for stay dates in period $t$ if inventory were unconstrained and matching were frictionless. Let $S_{t}$ denote effective supply: the number of listing-nights made available by hosts for stay dates in period $t$, defined as physical availability (active listing, unblocked night, not already reserved). Price ($P_{t}$) and platform visibility/ranking ($R_{t}$) are treated as separate covariates that influence both $D_{t}$ and $S_{t}$ but are not inventory in the same units as bookings. Let $B_{t}$ denote realized booked nights (the number of listing-nights that result in a completed reservation). Measurement issues affecting the construction of $S_{t}$, $B_{t}$, and demand proxies from external data sources are discussed in the Supplementary Material.
@@
    Available supply $S_{t}$: partially observed through scrapes, platform dashboards, or internal records. Treated here as “inventory” when scalar and “supply” when referring to the broader distributional object (composition by host type, price tier, location). Success metric: out-of-sample accuracy and measurement fidelity.
@@
The remainder of the paper is organized as follows. Section 2 reviews tourism demand forecasting since the Song et al. (2019) synthesis. Section 3 reviews platform accommodation markets, two-sided foundations, and the supply-side forecasting gap. Section 4 develops the coupling argument, using the censoring formulation to clarify when and why decoupled forecasts fail, and provides a simulation-based demonstration. Section 5 presents the research agenda with explicit identification assumptions and evaluation designs. Section 7 concludes.
2 Tourism demand forecasting since 2019
The post-2019 demand-forecasting literature is best read as incremental progress on a mature methodological core, plus a major diversification of data sources and evaluation settings. The most efficient framing is to position the Song et al. (2019) review as the methodological baseline and then note extensions along six frontiers.
Benchmark baselines and combination. A persistent finding is that simple methods remain competitive. Exponential smoothing, ARIMA, and naïve benchmarks continue to perform well against more complex alternatives (Athanasopoulos et al., 2011). Forecast combination has emerged as both a practical baseline and a research focus (Bates and Granger, 1969; Timmermann, 2006), with Song et al. (2021) demonstrating that bagging-based combination stabilizes accuracy in both normal and disrupted periods. Notably, combination gains arise precisely when individual models disagree, which is the expected condition when an omitted supply ceiling binds intermittently.
Deep learning. A common pattern combines decomposition with neural sequence models to mitigate overfitting (Zhang et al., 2021). Recurrent architectures are widely applied for longer-horizon forecasting (Salinas et al., 2020), while Transformer-based models have gained traction for capturing long-range dependencies (Lim et al., 2021). The appeal is the ability to capture nonlinear seasonality and abrupt level shifts without explicit structural specification; the limitation is sample requirements and limited interpretability. For coupled forecasting, the deeper concern is that flexible architectures can fit censored patterns during training without representing the censoring mechanism, making them vulnerable to the same regime-shift failures that affect simpler models (Section 4.5).
Alternative data and nowcasting. Search intensity proxies, web traffic, mobility signals, and transaction data now appear routinely as leading indicators (Sun et al., 2019; Li et al., 2021). A representative recent design uses point-of-sale card transactions to forecast daily tourism demand while integrating search trends, holidays, and weather (Grau-Escolano et al., 2026). The value of alternative data is most apparent during structural breaks, when traditional econometric covariates often weaken in predictive power. From the coupling perspective, these demand-intent proxies are especially valuable because they measure $D_{t}$ before the supply ceiling censors it into $B_{t}$; the gap between intent signals and realized bookings is itself a diagnostic for when coupling binds (Section 5).
Mixed-frequency modeling. Many fast signals arrive at daily or weekly frequency while core tourism targets remain monthly or quarterly. The MIDAS framework (Ghysels et al., 2004) and its deep-learning variants bridge this mismatch (Hu et al., 2025). This is especially relevant for coupled forecasting, because supply-side signals are inherently high-frequency while demand targets are often reported at lower aggregation levels.
Bayesian and ensemble approaches. Bayesian structural time series (Brodersen et al., 2015) and Bayesian global VAR models (Assaf et al., 2019) have been applied to tourism spillover analysis and arrival forecasting, with the latter demonstrating that cross-country demand linkages improve multi-step-ahead accuracy.
Competitions and crisis evaluation. The tourism forecasting competition tradition (Athanasopoulos et al., 2011) provided large-scale benchmarks. During COVID-19, the competition format encouraged scenario thinking and explicit judgmental adjustments (Li et al., 2022). The pandemic exposed a fundamental vulnerability of purely extrapolative methods and accelerated interest in methods that can incorporate external information and expert judgment. The coupling perspective sharpens this lesson: COVID-19 was simultaneously a demand shock (travel restrictions), a supply shock (host exit, regulatory tightening), and a compositional shock (redistribution across origins, lead times, and listing types). Methods that model only the demand margin cannot distinguish these channels.
Despite the breadth and depth of this literature, a consistent feature of tourism demand forecasting that targets arrivals or bookings is that demand is modeled as a standalone object, largely abstracting from endogenous capacity constraints relative to the revenue-management tradition. The demand forecasting toolkit is sophisticated and continues to improve, but its problem formulation is unchanged from the hotel era: predict arrivals given economic fundamentals and perhaps some leading indicators. This formulation becomes a substantive limitation in platform accommodation, where the supply environment is endogenous to the demand being forecasted.
@@ 3.2 Platform accommodation: pricing, supply dynamics, and regulation
The platform accommodation literature most relevant to coupled forecasting centers on pricing dynamics, supply dynamics, and regulatory intervention. On competitive impacts, Zervas et al. (2017) use difference-in-differences to estimate effects of Airbnb expansion on hotel performance. On pricing, heterogeneity by host type and professionalization means that host pricing rules are a behavioral channel coupling demand shocks into supply responses (Abrate and Viglia, 2022; Casamatta et al., 2022).
@@ 4 The coupling problem
The coupling operates through three distinct mechanisms.
4.1 Behavioral coupling: policies propagate across sides
Two-sided market theory establishes that platforms optimally set prices and rules by internalizing cross-group externalities (Rochet and Tirole, 2003). In practice, this means that policy changes nominally targeting one side inevitably affect the other. When a platform modifies cancellation rules, pricing recommendations, search ranking, or commission structures, the behavioral responses of both guests and hosts shift jointly, not independently.
Consider a well-documented example: the introduction of flexible cancellation or refund policies on short-term rental platforms. Jia et al. (2021) study Airbnb’s introduction of a 48-hour full-refund grace period and find that reservations increased, with effects decomposable into higher prices and higher occupancy. However, the same policy change propagates to the supply side. Lower commitment costs increase cancellation rates, altering the host experience. Hosts respond heterogeneously: professional operators absorb the volatility, while casual hosts, more sensitive to income variability, may reduce participation or exit.
The forecasting implication is direct. A demand model trained on pre-policy data will mis-forecast in the post-policy environment, not because demand fundamentals changed, but because the supply composition shifted in response to the same policy. Any platform intervention that alters incentives on one side will reshape participation and behavior on the other. Surge pricing creates feedback between demand signals and supply availability (Cachon et al., 2017). Changes to search ranking algorithms determine which listings are visible, effectively filtering the supply set. Behavioral coupling means that the marginal distribution of demand is not stable when supply-side conditions change, and vice versa. Critically, these platform interventions are chosen endogenously as functions of observed imbalance signals, not imposed exogenously; this endogeneity motivates the intervention-coupling problem below.
@@ 4.2 Informational coupling: the missing ceiling
This creates a specific informational problem for demand forecasting. For accommodation-focused targets in markets and periods where inventory binds, the observed outcome is censored by available supply. The censoring structure is stock-out censoring (Tobin, 1958): the analyst observes $B_{t}$ but not $D_{t}$ when $D_{t}>S_{t}$, and the censoring threshold $S_{t}$ is itself stochastic and endogenous. If the model does not include supply state variables, it is doing one of two things. Either it is implicitly assuming $S_{t}\to\infty$, in which case it will over-forecast in supply-tight markets. Or it is fitting to historically observed $B_{t}$ values that were already constrained by the prevailing supply environment, in which case it has learned the shape of the censoring mechanism without representing it, and it will fail when the supply regime changes.
The informational problem is compounded by a measurement problem. The demand side benefits from decades of convergence on well-defined targets: tourist arrivals, overnight stays, expenditures. No equivalent consensus exists for platform supply. Some studies count listings (Zervas et al., 2017), others measure spatial penetration (Quattrone et al., 2018), others estimate entry and exit flows (Fan et al., 2023; Kolleck and Teubner, 2025). The relevant supply forecast target is distributional: a characterization of available inventory by host type, location, price tier, and temporal availability pattern. Compositional time series methods (Aitchison, 1986; Katz et al., 2024) are designed for this class of target, but the forecasting literature has not yet framed the supply problem in compositional terms.
@@ 4.3 Intervention coupling: endogenous regime change
The third form of coupling arises because platforms and regulators intervene in accommodation markets, and these interventions are endogenous responses to observed supply-demand conditions that alter the structural relationship between the two sides.
@@
Platform design changes constitute a second, less observable class of intervention. Platform mechanisms such as search ranking, pricing recommendations, matching algorithms, and commission structures are periodically updated in response to observed market conditions. This creates a form of policy endogeneity familiar from the macroeconomics literature (Lucas, 1976): the historical relationship between supply and demand partly reflects past platform decisions, and those decisions are themselves functions of market state variables.
External shocks introduce a third variant. Pandemics, natural disasters, and exchange rate movements do not affect supply and demand symmetrically. They redistribute: some origin markets collapse while others grow; some listing types exit while others enter; lead-time distributions compress or stretch (Katz et al., 2025). These are compositional changes: changes in shape, not merely level. Standard structural break tools that detect level shifts in scalar time series miss the directionality and dimensionality of the disruption. Katz (2026) proposes a Bayesian Dirichlet ARMA model with a directional-shift intervention mechanism for this class of problem.
For analytical clarity, it is useful to distinguish three types of structural breaks in coupled systems: (i) demand shocks that shift $D_{t}$ while leaving $S_{t}$ initially unchanged; (ii) supply shocks that shift $S_{t}$ while leaving $D_{t}$ initially unchanged; and (iii) platform intervention shocks that simultaneously alter the coupling relationship between $D_{t}$ and $S_{t}$. Standard break detection methods are best suited to type (i) and (ii) breaks. Type (iii) breaks, which manifest as changes in coupling parameters rather than in either marginal series, require multivariate extensions discussed in the research agenda.

### 19 Apr 2026 v2

# Coupled Supply and Demand Forecasting in Platform Accommodation Markets Harrison E. Katz

@@ Abstract
Tourism demand forecasting is methodologically mature, but it typically treats accommodation supply as fixed or exogenous. In platform-mediated short-term rentals, supply is elastic, decision-driven, and co-evolves with demand through pricing, information design, and interventions. Realized booked nights cannot exceed either latent demand or available supply, so booking models that ignore supply learn a regime-specific ceiling and become fragile under policy changes and supply shocks. This narrated review synthesizes work from tourism forecasting, revenue management, two-sided market economics, and Bayesian time-series methods; develops a three-part coupling framework spanning intervention, behavioral, and informational channels; and illustrates the identification failure with a toy simulation. I conclude with a focused research agenda for jointly forecasting supply, demand, and their compositions.
@@ 1 Introduction
The crucial shift is that “capacity” is no longer slow-moving and near-fixed (as in most hotel markets), but elastic at multiple margins: host entry and exit, listing activation and deactivation, and day-level availability decisions. Empirically, peer-to-peer accommodation supply is more flexible than hotel capacity, and the welfare gains from peer entry are largest in locations and periods where hotel capacity binds (Farronato and Fradkin, 2022). Host entry, exit, and calendar activity also show that effective platform inventory varies at multiple margins (Fan et al., 2023; Wang et al., 2024). From a theory standpoint, two-sided market economics provides the backbone for why forecasting demand alone is incomplete in such settings: platforms internalize cross-side externalities via pricing, matching, governance, and information design (Rochet and Tirole, 2003; Weyl, 2010; Armstrong, 2006).
@@
The paper makes three contributions. First, it develops a coupling framework that distinguishes intervention coupling (rule or regime changes), behavioral coupling (host and guest responses to those changes), and informational coupling (the fact that bookings are observed under a moving and only partly observed supply ceiling). Second, it uses a stylized simulation to isolate one identification failure: after a supply shock, demand-only forecasters systematically over-predict because they have learned a regime-specific ceiling from the training period. Third, it outlines a five-problem research agenda whose resolution would constitute the core of a coupled forecasting research program for platform markets. A detailed survey of the methodological building blocks for coupled forecasting is provided in the Supplementary Material.
@@
Let $D_{t}$ denote latent demand intent: the number of listing-nights that guests would book for stay dates in period $t$ if inventory were unconstrained and matching were frictionless. Let $S_{t}$ denote effective supply: the number of listing-nights made available by hosts for stay dates in period $t$, defined as physical availability (active listing, unblocked night, not already reserved). Price ($P_{t}$) and platform visibility/ranking ($R_{t}$) are treated as separate covariates that influence both $D_{t}$ and $S_{t}$ but are not inventory in the same units as bookings. Let $B_{t}$ denote realized booked nights (the number of listing-nights that result in a completed reservation). Measurement issues affecting the construction of $S_{t}$, $B_{t}$, and demand proxies from external data sources are summarized below and expanded in the Supplementary Material.
@@
This segment-level formulation also clarifies scope. The ceiling logic is strongest at the listing level, within narrow segments of close substitutes, and for short-run city forecasting after identifiable supply shocks. At broader aggregation or longer horizons, substitution across neighborhoods or property types, price adjustment, and host entry or re-entry can attenuate the constraint. The single-segment benchmark is therefore an analytic device, not a claim that every market-level series is equally ceiling-dominated.
@@
    Available supply $S_{t}$: partially observed and often noisy. Public calendar scrapes recover an imperfect proxy because blocked and booked nights are confounded; platform-internal aggregates are cleaner but usually unavailable to external researchers. When observables are listing-level (views, wishlists, dashboard states), market-level use requires explicit aggregation across listings or segments. Treated here as “inventory” when scalar and “supply” when referring to the broader distributional object (composition by host type, price tier, location). Success metric: out-of-sample accuracy and measurement fidelity.
Measurement, observability, and aggregation
Three empirical distinctions are important for the rest of the paper. First, public calendar scrapes and platform-internal records are different data environments. Public scrapes can approximate availability and occupancy changes, but they cannot cleanly distinguish blocked nights from booked nights. Internal records can do so directly. Second, several proposed observables, including listing views, wishlists, and dashboard metrics, are naturally listing-level objects. They become city-level or segment-level variables only if the platform aggregates them across listings. Third, scalar demand-intent proxies and distributional demand-intent proxies should not be conflated. Google Trends, card transactions, online reviews, and related public indicators can proxy aggregate demand intensity, but they do not recover booking lead-time distributions. The compositional version of the search-to-booking gap therefore requires internal search logs or equivalently granular search data.
Table 1: Data environments relevant for coupled forecasting in platform accommodation.
Data environment Examples What it can identify Main limitation
Public scrape data Repeated calendar snapshots; listing pages; vendor panels Approximate availability, listing counts, occupancy changes, posted prices Blocked and booked nights look identical; minimum-stay rules and coverage gaps create measurement error
Public aggregate demand proxies Google Trends, card transactions, online review volume and valence Scalar demand-intent signals and nowcasting covariates Coarse aggregation; no lead-time composition; imperfect mapping to booking intent
Platform-internal aggregated data Aggregated searches, views, wishlists, booked nights, exact calendar states Cleaner measures of $B_{t}$ and $S_{t}$; lead-time-specific conversion gaps; market-level aggregation from listing data Usually unavailable to external researchers
Administrative and regulatory data Registration rolls, permit records, enforcement dates Intervention timing, legal supply restrictions, external validation for supply shocks Incomplete coverage of informal supply; often lagged relative to platform state
@@
The remainder of the paper is organized as follows. Section 2 reviews recent developments in tourism demand forecasting after the synthesis of Song et al. (2019). Section 3 reviews platform accommodation markets, two-sided foundations, and the supply-side forecasting gap. Section 4 develops the coupling argument, clarifies the three coupling layers, and provides a stylized simulation-based demonstration. Section 5 presents the research agenda with explicit identification assumptions, data environments, and evaluation designs. Section 6 discusses practitioner implications. Section 7 states the boundary conditions and limitations of the framework. Section 8 concludes.
2 Recent developments in tourism demand forecasting
The recent tourism demand forecasting literature is better understood as the continuation of a mature forecasting toolkit under new data and new stress tests, rather than as a sequence of disconnected technical novelties. A useful recent stock-taking in Tourism Management is Song et al. (2023), which revisits the field’s theoretical and empirical development. For the present paper, three developments matter most: renewed attention to stabilization and combination, richer and faster data, and greater sensitivity to regime instability.
First, the literature continues to find that strong simple baselines remain competitive, which is why combination and stabilization have become central. Athanasopoulos et al. (2011) show why such baselines remain essential in tourism forecasting: automated time-series methods perform well, and naïve annual forecasts are hard to beat. Forecast combination has therefore remained both a practical baseline and an active research frontier (Bates and Granger, 1969; Timmermann, 2006), with Song et al. (2021) showing that bagging-based combination can stabilize accuracy in both normal and disrupted periods. Recent Tourism Management work refines this point by showing that decomposition choices materially affect bagging performance (Liu et al., 2023). Related work on decomposition and nonlinear pattern recognition can be read in the same light: recurrent networks, Transformers, and compound pattern-recognition systems are attractive because they can absorb irregular seasonality and abrupt shifts (Zhang et al., 2021; Salinas et al., 2020; Lim et al., 2021; Hu et al., 2025b). In a coupled accommodation market, however, even flexible forecasting architectures can be misled if the target series is censored by an omitted or poorly measured supply ceiling.
Second, the literature increasingly expands the information set beyond conventional macroeconomic covariates. Search intensity, internet-activity measures, card-transaction data, and online reviews now appear routinely as leading indicators in tourism forecasting (Sun et al., 2019; Li et al., 2021; Hu et al., 2022; Grau-Escolano et al., 2026). Recent Tourism Management studies show that tourist-generated online reviews can improve destination forecasting (Hu et al., 2022). Mixed-frequency and multi-source designs that combine high-frequency digital traces with lower-frequency aggregate information can also materially improve forecasts during COVID-era recovery (Wu et al., 2023; Hu et al., 2025a). At the same time, recent critical assessment warns that search-query applications are highly sensitive to query design, preprocessing, and model specification (Mikulić and Baumgärtner, 2025). From the perspective of this paper, the key point is not simply that these variables can improve accuracy. Rather, they widen the information set around evolving demand conditions. Some of them, especially search and web-traffic measures, may proxy demand interest earlier and at higher frequency than bookings do, while others help characterize shifts in market conditions that are only imperfectly reflected in booking data alone. This is especially useful when observed bookings may be censored by a moving supply ceiling.
Third, the pandemic made regime instability a first-order forecasting problem. The tourism forecasting competition tradition already emphasized disciplined benchmarking (Athanasopoulos et al., 2011), but COVID-19 accelerated scenario thinking, explicit judgmental adjustment, and the use of external information (Song et al., 2022). This matters for the coupling argument because the pandemic was not a pure demand shock. It simultaneously changed travel intent, host participation, regulatory conditions, and the composition of travel across origins, lead times, and listing types. The recent literature therefore illustrates a broader point: progress in tourism demand forecasting has been substantial, but the forecasting object is still usually framed as demand evolving under an implicit or only slowly varying supply environment.
That formulation remains appropriate when capacity is approximately fixed or the supply environment changes slowly relative to the forecast horizon. It becomes a substantive limitation in platform accommodation, where supply is elastic, decision-driven, and endogenous to the demand being forecast.
@@ 3.2 Platform accommodation: pricing, supply dynamics, and regulation
The platform accommodation literature most relevant to coupled forecasting centers on pricing dynamics, supply dynamics, and regulatory intervention. On competitive impacts, Zervas et al. (2017) use difference-in-differences to estimate effects of Airbnb expansion on hotel performance. On pricing, host heterogeneity and professionalization shape how Airbnb listings are priced across market states (Abrate et al., 2022; Casamatta et al., 2022). This matters for the coupling argument because pricing is one channel through which demand shocks can feed back into host-side participation and availability. Recent Tourism Management evidence from European Airbnb markets during the pandemic also shows that response-stringency measures and demand shifts propagated into listing prices, with substantial differences between commercial and private hosts (Milone et al., 2023).
@@ 4 The coupling problem
The coupling operates through three distinct mechanisms. Analytically, intervention coupling is the trigger, behavioral coupling is the propagation channel, and informational coupling is the observability problem. In practice these mechanisms interact recursively, but separating their roles clarifies why different forecasting failures arise.
Intervention coupling
Rule or regime changes
(regulation, ranking, commissions, policy redesign)
$\rightarrow$ Behavioral coupling
Host and guest responses
(entry/exit, pricing, availability, booking and cancellation behavior)
$\rightarrow$ Informational coupling
Observed bookings under a moving, partly observed ceiling
(search-to-booking divergence, stock-out censoring)
Forecasting implication: decoupled models transfer parameters across regimes, miss cross-side propagation, and misread booked nights as unconstrained demand.
Figure 1: Conceptual ordering of the coupling framework. Intervention coupling changes the rules of the market, behavioral coupling transmits the change through host and guest responses, and informational coupling determines what the forecaster observes in bookings. The mechanisms interact recursively in practice, but the ordering clarifies their analytical roles.
4.1 Behavioral coupling: cross-side propagation
Behavioral coupling refers to the host-side and guest-side responses that transmit a shock, rule change, or market signal across sides. The intervention itself is analytically distinct. A regulator or platform may change a rule, but behavioral coupling is the way hosts and guests respond through entry and exit, pricing, availability, search, booking, and cancellation behavior.
Two-sided market theory provides the mechanism. Because participation on one side depends on expected participation on the other, a shock that initially lands on one side rarely stays there (Rochet and Tirole, 2003). Consider a well-documented example: Airbnb’s introduction of a 48-hour full-refund grace period. Jia et al. (2021) show that Airbnb’s pro-guest cancellation-policy changes shifted both demand and supply on the platform, with heterogeneous effects across host types. That makes the case especially useful here: a guest-facing rule can propagate through prices, occupancy, and host participation rather than remaining confined to one side of the market.
The forecasting implication is that cross-side propagation is endogenous to the market state. A demand model trained before such responses occur will mis-forecast afterward, not only because the rule changed, but because the host and guest responses reshape the participation margin on both sides. Host exit and entry, calendar activation and blocking, price resetting, and cancellation responses are therefore better viewed as behavioral manifestations of coupling. By contrast, rule changes such as commission adjustments, ranking changes, or cancellation-policy redesign are intervention variables discussed in Section 4.3.
@@ 4.2 Informational coupling: the missing ceiling
This creates a specific informational problem for demand forecasting. For accommodation-focused targets in markets and periods where inventory binds, the observed outcome is censored by available supply. The censoring structure is akin to a censored or unconstrained-demand problem: the analyst observes $B_{t}$ but not $D_{t}$ when $D_{t}>S_{t}$, and the censoring threshold $S_{t}$ is itself stochastic and endogenous (Tobin, 1958; Talluri and van Ryzin, 2004). If the model does not include supply state variables, it is doing one of two things. Either it is implicitly assuming $S_{t}\to\infty$, in which case it will over-forecast in supply-tight markets. Or it is fitting to historically observed $B_{t}$ values that were already constrained by the prevailing supply environment, in which case it has learned the shape of the censoring mechanism without representing it, and it will fail when the supply regime changes.
The informational problem is compounded by a measurement problem. The demand side benefits from decades of convergence on well-defined targets: tourist arrivals, overnight stays, expenditures. No equivalent consensus exists for platform supply. Some studies count listings (Zervas et al., 2017), others measure spatial penetration (Quattrone et al., 2018), others estimate entry and exit flows (Fan et al., 2023; Kolleck and Teubner, 2025). The relevant supply forecast target is distributional: a characterization of available inventory by host type, location, price tier, and temporal availability pattern. Compositional data methods provide a natural starting point for this class of target, because they are built for constrained share vectors rather than unconstrained scalar series (Aitchison, 1986; Katz et al., 2024). The forecasting literature has not yet framed the supply problem explicitly in those terms.
@@ 4.3 Intervention coupling: endogenous regime change
Intervention coupling concerns rule or regime changes that alter the mapping between supply, demand, and observed bookings. The defining feature is not that behavior changes. Behavior also changes under many ordinary market shocks. The defining feature is that a platform or regulator changes the institutional environment itself, often in response to the current state of the market.
@@
Platform design changes constitute a second, less observable class of intervention. Search ranking, pricing recommendations, matching rules, fee structures, and cancellation-policy redesigns are periodically updated in response to observed market conditions. This creates a form of policy endogeneity familiar from the macroeconomics literature (Jr, 1976): the historical relationship between supply and demand partly reflects past platform decisions, and those decisions are themselves functions of market state variables.
External shocks introduce a third variant when they force a regime change in the coupling relationship itself. The pandemic, in particular, redistributed demand across origins, listing types, and lead-time profiles rather than merely shifting aggregate levels (Katz et al., 2025; Yang et al., 2021). These are compositional changes, not merely level changes. Standard structural-break tools that detect level shifts in scalar time series miss this directionality and dimensionality. One recent proposal is Katz (2026), which develops a Bayesian Dirichlet ARMA model with a directional-shift intervention mechanism for compositional structural breaks.
For analytical clarity, it is useful to distinguish three types of structural breaks in coupled systems: (i) demand shocks that shift $D_{t}$ while leaving $S_{t}$ initially unchanged; (ii) supply shocks that shift $S_{t}$ while leaving $D_{t}$ initially unchanged; and (iii) platform intervention shocks that alter the coupling relationship between $D_{t}$ and $S_{t}$. Standard break-detection methods are best suited to type (i) and type (ii) breaks. Type (iii) breaks, which manifest as changes in coupling parameters rather than in either marginal series, require the multivariate extensions discussed in the research agenda.